# 3D-Properties: Identifying Challenges in DPO and Charting a Path Forward

### v1

@@ 6 Concluding Remarks and Limitations
In this study, we conducted a comprehensive theoretical analysis to elucidate why the widely used alignment algorithm DPO, does not perform as well as the RLHF-PPO algorithm. The principal challenge identified in DPO is summarized as 3D-properties. We substantiated our theoretical framework through experimental results obtained from both a toy model and real Large Language Models (LLMs) in practical applications, including mathematical reasoning and instruction following. Additionally, we assessed the effectiveness of specific regularization techniques. Furthermore, by contrasting DPO training with reward model training, we highlighted the inherent instability of DPO. We hope this work could offer research directions to narrow the gap between reward-free preference learning methods and reward-based ones. One limitation of this study is that we have not yet tested our theory on the recently proposed token-level reward DPO algorithm [37]. We plan to address this gap in future research.
@@ Appendix A Detailed Background and Related Works
Large language models (LLMs) are profoundly transforming the way we work and live. Performing a three-stage process is the default practice for training LLMs: Pretraining, Supervised Fine-Tuning (SFT), and Reinforcement Learning from Human Feedback (RLHF). The roles of Pretraining and SFT are broadly understood: Pretraining encodes knowledge and SFT aligns question-answer formats. Relatively speaking, the understanding of RLHF is relatively insufficient. Specifically, Direct Preference Optimization (DPO) and its variants, as reward-model-free algorithms, have garnered significant attention due to their elegant mathematical form and relatively low resource requirements [23, 22, 13, 29]. However, it has also sparked considerable debate because of its unstable performance in practical applications [17, 31].
@@
The key inspiration for the DPO algorithm [23] is a closed-form solution to the RL step in RLHF, and thus an equivalent solution to the optimal policy for RLHF objective. The original DPO work is an off-policy learning algorithm for it relies on an extra preference dataset (Helpful-and-Harmless [6]), where the preference pairs are not generated by the policy LLM itself. On the other hand, there are a bunch of on-policy learning algorithms developed, where the preference responses are sampled from the policy model. [13] proposed the on-policy version of DPO. In on-policy DPO, all responses are sampled in a batch-wise way. A natural trade-off between them is the iterative DPO introduced by [29, 31]. The algorithm begins by initializing with an additional preference dataset, then iteratively trains a policy using DPO, collects response pairs through exploration policies, obtains preference signals from human or AI labelers, and updates the dataset with the newly labeled data.
@@ A.2 Insights into DPO
Though the reward-free algorithms are favored due to their lower computational overhead, if they can achieve on-performance with state-of-art reward-based methods such as RLHF-PPO sparked a lot of discussions. [18] proves that the absence of a reward model in DPO constrains its ability to sample preference pairs from the optimal policy. [31] show that DPO may have fundamental limitations that its optimal solution is a superset of the optimal solution of the PPO algorithm. This work also reports the empirical results that the performance of DPO is affected by the distribution shift between the model outputs and the preference dataset. [12] discusses the limitations of DPO from the perspective of gradient numerical stability, and conducted experiments to preliminarily verify it. However, they did not conduct experiments on real LLM and illustrate the correlation.
A.3 Other reward-free alignment algorithms
A major limitation of the DPO objective is its reliance on the Bradley-Terry model to convert pairwise preferences into point-wise rewards. To overcome this, [4] introduced $\Phi$-preference optimization ($\Phi\text{PO}$), where DPO is a special case of it that $\Phi(P)=\log\frac{P}{1-P}$. Identity-preference optimization (IPO) is a variant that replaces the $\Phi$-function by an identity mapping function $\Phi(P)=P$.
Different from the DPO or IPO, the core idea of Sequence Likelihood Calibration (SLiC) [36] is to calibrate the likelihood of ranked sequences sampled from the policy being trained. The SLiC loss function can be decomposed into two parts: the rank function to guarantee that the difference between $\log\pi_{\theta}(a^{+}|x)$ and $\log\pi_{\theta}(a^{-}|x)$ is greater than $\delta$ under the current policy $\pi_{\theta}$, and the cross-entropy regularizer that to encourage the model to stay close to the SFT model.
There are some other variants that tries to improve DPO, such as KTO [11], NCA [8], ODPO (DPO with an offset) [2]. KTO uses a Kahneman-Tversky model of human utility and proposes a method that directly maximizes the utility of generations instead of maximizing the log-likelihood of preferences. NCA leverages Noise Contrastive Estimation (NCE) to bridge the gap in handling reward datasets explicitly annotated with scalar evaluations. ODPO does not treat every preference pair equally during fine-tuning and requires the difference between the likelihood of the preferred and dispreferred response to be greater than an offset value.
Appendix B Theoretical Foundation
@@ B.1 Fundamental Limitation in Vanilla DPO
Here we revisit the theoretical findings in Section 3.1. The loss function for vanilla DPO is given by
@@
$\alpha:=\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\right)^{\beta},\quad\pi^{+}:=\pi(a^{+}|x),\quad\pi^{-}:=\pi(a^{-}|x),$
and
$z=\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}.$
@@
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}\to-\alpha\beta(\pi^{+})^{-\beta-1}(\pi^{-})^{\beta},$
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}\to\alpha\beta(\pi^{+})^{-\beta}(\pi^{-})^{\beta-1}.$
@@
$\pi^{+}=\prod_{i}\pi^{+}_{i}=\pi_{-i}^{+}\cdot\pi_{i}^{+},$
$\pi^{-}=\prod_{i}\pi^{-}_{i}=\pi_{-i}^{-}\cdot\pi_{i}^{-}.$
Here we have
$\frac{\partial\pi}{\partial\pi_{i}}=\pi_{-i}.$
Consider a softmax function:
$s_{i}=\frac{e^{z_{i}}}{\sum_{j}e^{z_{j}}}.$
The corresponding gradients are
@@
$C(\pi^{+},\pi^{-}):=\alpha\beta^{+}(\pi^{+})^{-\beta^{+}}(\pi^{-})^{\beta^{-}}.$
Therefore, considering the current selection probability $\pi^{+}_{i}$ of the $i$-th token for the chosen sequence, let the sampled token’s index be $c$. The logit corresponding to this token $c$ is denoted as $x^{+}_{i,c}$, then we have
@@
Therefore, considering the current selection probability $\pi^{-}_{i}$ of the $i$-th token for the rejected sequence, let the sampled token’s index be $c$. The logit corresponding to this token $c$ is denoted as $x^{-}_{i,c}$, then we have
@@ B.2.1 Flexible $\beta$-DPO
$\alpha:=\frac{\pi_{0}(a^{+}|x)^{\beta^{+}}}{\pi_{0}(a^{-}|x)^{\beta^{-}}},\quad\pi^{+}:=\pi(a^{+}|x),\quad\pi^{-}:=\pi(a^{-}|x),$
and
$z=\frac{\pi_{\theta}(a^{-}|x)^{\beta^{-}}}{\pi_{\theta}(a^{+}|x)^{\beta^{+}}}.$
@@
$\frac{\partial\ell^{\text{DPO'}}}{\partial\pi^{+}}\to-\alpha\beta^{+}(\pi^{+})^{-\beta^{+}-1}(\pi^{-})^{\beta^{-}},$
$\frac{\partial\ell^{\text{DPO'}}}{\partial\pi^{-}}\to\alpha\beta^{-}(\pi^{+})^{-\beta^{+}}(\pi^{-})^{\beta^{-}-1}.$
Using different $\beta^{+}$ and $\beta^{-}$ can only alleviate 3D-properties mentioned in Section 3.
@@ B.2.2 SFT Loss Regularization
As $\pi^{-}\to 0$, the gradient for the chosen action $\frac{\partial\ell^{\text{SFT-DPO}}}{\partial\pi^{+}}\to-\gamma\frac{1}{\pi^{+}}\neq 0$, which means the likelihood of chosen responses can be continually optimized.
@@ B.3.1 Identity-preference Optimization (IPO)
Figure 5: IPO performance on toy model.
Figure 6: IPO Gradient and Log probability.
@@ B.3.2 Sequence Likelihood Calibration (SLiC)
Figure 7: SLiC performance on toy model.
Figure 8: SLiC Gradient and Log probability.
@@
These variants all alleviate the 3D-properties problem of DPO, so the results on mathematical reasoning are partly improved (see Table 10).
@@
The invariants we tested do not perform well uniformly on all the tasks. For example, on instruction following tasks like poem generation, vanilla DPO outperforms SLiC and IPO. One hypothesis is their solution forms diverge significantly from the Bradley-Terry model, leading to a loss of generalization in preference learning.
@@ Appendix C Dataset Description
Setting MATH SuperCLUE Poem Slogan COMMON
train set 5826 - 93269 13592 400000
test set 2000 1072 1000 1000 -
@@
Table 7 shows the amount of data in each dataset. MATH and SuperCLUE are open-source datasets, while Poem, Slogan, and COMMON are in-house self-built datasets. We provide part of the in-house datasets in the supplementary material to clarify the format and the content. SuperCLUE is only for cross-dataset testing, and COMMON is only used for the comparison between the reward model training and DPO training in Section 5.3.

### v2

@@ 5 Conclusions
In this study, we conducted a comprehensive theoretical analysis to elucidate why DPO does not perform as well as RM-based alignment algorithm. The principal challenge identified in DPO is summarized as 3D-properties. We substantiated our theoretical framework through experimental results obtained from both a toy model and real LLMs in practical applications, including mathematical reasoning and instruction following. Additionally, we assessed the effectiveness of specific regularization techniques. Furthermore, by contrasting DPO training with RM training, we highlighted the inherent instability of DPO. We hope this work could offer research directions to narrow the gap between RM-free preference learning methods and RM-based ones. We leave the discussion of limiation to Appendix D.3.
@@ Appendix A Detailed Background and Related Works
Large language models (LLMs) are profoundly transforming the way we work and live. Performing a three-stage process is the default practice for training LLMs: Pretraining, Supervised Fine-Tuning (SFT), and Reinforcement Learning from Human Feedback (RLHF). The roles of Pretraining and SFT are broadly understood: Pretraining encodes knowledge and SFT aligns question-answer formats. Relatively speaking, the understanding of RLHF is relatively insufficient. Specifically, Direct Preference Optimization (DPO) and its variants, as reward-model-free algorithms, have garnered significant attention due to their elegant mathematical form and relatively low resource requirements (Rafailov et al., 2024; Pal et al., 2024; Guo et al., 2024; Xiong et al., 2023). However, it has also sparked considerable debate because of its unstable performance in practical applications (Li et al., 2023; Xu et al., 2024a).
@@
The key inspiration for the DPO algorithm (Rafailov et al., 2024) is a closed-form solution to the RL step in RLHF, and thus an equivalent solution to the optimal policy for RLHF objective. The original DPO work is an off-policy learning algorithm for it relies on an extra preference dataset (Helpful-and-Harmless (Bai et al., 2022)), where the preference pairs are not generated by the policy LLM itself. On the other hand, there are a bunch of on-policy learning algorithms developed, where the preference responses are sampled from the policy model. Guo et al. (2024) proposed the on-policy version of DPO. In on-policy DPO, all responses are sampled in a batch-wise way. A natural trade-off between them is the iterative DPO introduced by Xiong et al. (2023); Xu et al. (2024a). The algorithm begins by initializing with an additional preference dataset, then iteratively trains a policy using DPO, collects response pairs through exploration policies, obtains preference signals from human or AI labelers, and updates the dataset with the newly labeled data.
@@ A.2 Insights into DPO
Though the RM-free algorithms are favored due to their lower computational overhead, if they can achieve on-performance with state-of-art RM-based methods such as RLHF-PPO sparked a lot of discussions. Liu et al. (2023) proves that the absence of RM in DPO constrains its ability to sample preference pairs from the optimal policy. Xu et al. (2024a) show that DPO may have fundamental limitations that its optimal solution is a superset of the optimal solution of the PPO algorithm. This work also reports the empirical results that the performance of DPO is affected by the distribution shift between the model outputs and the preference dataset. Feng et al. (2024) discusses the limitations of DPO from the perspective of gradient numerical stability, and conducted experiments to preliminarily verify it. However, they did not conduct experiments on real LLM and illustrate the correlation.
A.3 Other RM-free alignment algorithms
A major limitation of the DPO objective is its reliance on the Bradley-Terry model to convert pairwise preferences into point-wise rewards. To overcome this, Azar et al. (2024) introduced $\Phi$-preference optimization ($\Phi\text{PO}$), where DPO is a special case of it that $\Phi(P)=\log\frac{P}{1-P}$. Identity-preference optimization (IPO) is a variant that replaces the $\Phi$-function by an identity mapping function $\Phi(P)=P$.
Different from the DPO or IPO, the core idea of Sequence Likelihood Calibration (SLiC) (Zhao et al., 2023) is to calibrate the likelihood of ranked sequences sampled from the policy being trained. The SLiC loss function can be decomposed into two parts: the rank function to guarantee that the difference between $\log\pi_{\theta}(a^{+}|x)$ and $\log\pi_{\theta}(a^{-}|x)$ is greater than $\delta$ under the current policy $\pi_{\theta}$, and the cross-entropy regularizer that to encourage the model to stay close to the SFT model.
There are some other variants that tries to improve DPO, such as KTO (Ethayarajh et al., 2024), NCA (Chen et al., 2024), ODPO (DPO with an offset) (Amini et al., 2024). KTO uses a Kahneman-Tversky model of human utility and proposes a method that directly maximizes the utility of generations instead of maximizing the log-likelihood of preferences. NCA leverages Noise Contrastive Estimation (NCE) to bridge the gap in handling reward datasets explicitly annotated with scalar evaluations. ODPO does not treat every preference pair equally during fine-tuning and requires the difference between the likelihood of the preferred and dispreferred response to be greater than an offset value.
Appendix B Theoretical Foundations
@@ B.1 Fundamental Limitation in Vanilla DPO
Here we revisit the theoretical findings in Section 3.1 in detail. The loss function for vanilla DPO is given by
@@
$\alpha:=\left(\frac{\pi_{0}(a^{+}|x)}{\pi_{0}(a^{-}|x)}\right)^{\beta},\quad\pi^{+}:=\pi_{\theta}(a^{+}|x),\quad\pi^{-}:=\pi_{\theta}(a^{-}|x),\quad z:=\frac{\pi_{\theta}(a^{-}|x)}{\pi_{\theta}(a^{+}|x)}.$
@@
$\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{+}}\to-\alpha\beta(\pi^{+})^{-\beta-1}(\pi^{-})^{\beta},\quad\frac{\partial\ell^{\text{DPO}}}{\partial\pi^{-}}\to\alpha\beta(\pi^{+})^{-\beta}(\pi^{-})^{\beta-1}.$
@@
$\pi^{+}=\prod_{i}\pi^{+}_{i}=\pi_{-i}^{+}\cdot\pi_{i}^{+},\quad\pi^{-}=\prod_{i}\pi^{-}_{i}=\pi_{-i}^{-}\cdot\pi_{i}^{-},$
where we have $\frac{\partial\pi}{\partial\pi_{i}}=\pi_{-i}$. Consider a softmax function, $s_{i}=\frac{e^{z_{i}}}{\sum_{j}e^{z_{j}}}$, the corresponding gradients are
@@
$C(\pi^{+},\pi^{-}):=\alpha\beta^{+}(\pi^{+})^{-\beta^{+}}(\pi^{-})^{\beta^{-}},$
then, consider the current selection probability $\pi^{+}_{i}$ of the $i$-th token for the chosen sequence, let the sampled token’s index be $c$. The logit corresponding to this token $c$ is denoted as $x^{+}_{i,c}$, then we have
@@
Similarly, consider the current selection probability $\pi^{-}_{i}$ of the $i$-th token for the rejected sequence, let the sampled token’s index be $c$. The logit corresponding to this token $c$ is denoted as $x^{-}_{i,c}$, then we have
@@ B.2.1 Flexible $\beta$-DPO
$\alpha:=\frac{\pi_{0}(a^{+}|x)^{\beta^{+}}}{\pi_{0}(a^{-}|x)^{\beta^{-}}},\quad\pi^{+}:=\pi(a^{+}|x),\quad\pi^{-}:=\pi(a^{-}|x),\quad z:=\frac{\pi_{\theta}(a^{-}|x)^{\beta^{-}}}{\pi_{\theta}(a^{+}|x)^{\beta^{+}}}.$
@@
$\frac{\partial\ell^{\text{DPO'}}}{\partial\pi^{+}}\to-\alpha\beta^{+}(\pi^{+})^{-\beta^{+}-1}(\pi^{-})^{\beta^{-}},\quad\frac{\partial\ell^{\text{DPO'}}}{\partial\pi^{-}}\to\alpha\beta^{-}(\pi^{+})^{-\beta^{+}}(\pi^{-})^{\beta^{-}-1}.$
As shown by the expressions for the gradients of the DPO’ objective with respect to the chosen and rejected response likelihoods, the magnitude of the gradients is controlled by the parameters $\beta^{+}$ and $\beta^{-}$. In particular, increasing $\beta^{+}$ strengthens the gradient for the chosen response, $\pi^{+}$, while reducing $\beta^{-}$ dampens the gradient for the rejected response, $\pi^{-}$, as it approaches zero. This gradient behavior suggests that adjusting these parameters can effectively mitigate the 3D-properties discussed in Section 3. Specifically, a large $\beta^{+}$ ensures that the likelihood of the chosen responses remains sufficiently reinforced, while a small $\beta^{-}$ prevents the likelihood of rejected responses from decreasing too rapidly, which would otherwise lead to the instability and degradation described earlier.
By fine-tuning $\beta^{+}$ and $\beta^{-}$, it becomes possible to control the interaction between the gradients of the chosen and rejected responses, reducing the drastic drop in rejected response likelihood, the degradation into response suppression, and the dispersion effect on unseen responses. This strategy thus offers a potential solution for improving the stability and performance of DPO by reducing the severity of the 3D-properties.
@@ B.2.2 SFT Loss Regularization
As $\pi^{-}\to 0$, the gradient for the chosen action $\frac{\partial\ell^{\text{SFT-DPO}}}{\partial\pi^{+}}\to-\gamma\frac{1}{\pi^{+}}\neq 0$, meaning that the likelihood of the chosen responses can continue to be optimized. This behavior is significant, as it indicates that even as the likelihood of rejected responses $\pi^{-}$ approaches zero, the chosen response $\pi^{+}$ can still be improved. This is a key advantage of SFT-DPO over the vanilla DPO, which suffers from gradient vanishing for $\pi^{+}$ when $\pi^{-}\to 0$. The negative impact of 3D-properties is thus reduced, allowing for more stable and effective optimization in the long run.
@@ B.3.2 Sequence Likelihood Calibration (SLiC)
B.3.3 Simple Preference Optimization (SimPO)
In SimPO, the loss function is written by,
$\ell^{\text{SimPO}}(\theta)=-\sum_{x,a^{+},a^{-}}\left[\log\sigma\left(\frac{\beta}{|a^{+}|}\log\pi_{\theta}(a^{+}|x)-\frac{\beta}{|a^{-}|}\log\pi_{\theta}(a^{-}|x)-\gamma\right)\right],$ (13)
where ||$\cdot$|| ⋅ | represents the length of the generated response and $\gamma$ is a target reward margin term. Similarly, we directly give the gradients:
$\frac{\partial\ell^{\text{SimPO}}}{\partial\pi_{\theta}(a^{+}|x)}=-\frac{\beta}{|a^{+}|\cdot\pi_{\theta}(a^{+}|x)}\sigma\left(-\left(\frac{\beta}{|a^{+}|}\log\pi_{\theta}(a^{+}|x)-\frac{\beta}{|a^{-}|}\log\pi_{\theta}(a^{-}|x)-\gamma\right)\right),$ (14)
$\frac{\partial\ell^{\text{SimPO}}}{\partial\pi_{\theta}(a^{-}|x)}=\frac{\beta}{|a^{-}|\cdot\pi_{\theta}(a^{-}|x)}\sigma\left(-\left(\frac{\beta}{|a^{+}|}\log\pi_{\theta}(a^{+}|x)-\frac{\beta}{|a^{-}|}\log\pi_{\theta}(a^{-}|x)-\gamma\right)\right).$ (15)
The ratio of the gradient is
$\frac{\partial\ell^{\text{SimPO}}}{\partial\pi^{-}}/\frac{\partial\ell^{\text{SimPO}}}{\partial\pi^{+}}=-\frac{|a^{+}|\cdot\pi_{\theta}(a^{+}|x)}{|a^{-}|\cdot\pi_{\theta}(a^{-}|x)},$ (16)
which indicates that as $\pi^{+}$ increases and $\pi^{-}$ decreases, the gradient with respect to $\pi^{-}$ grows faster and leads to a rapid drop in the likelihood of the rejected response. However, different from DPO, as $\pi^{-}\rightarrow 0$, we have $\partial\ell^{\text{SimPO}}/\partial\pi^{+}\rightarrow 0$. As for $\partial\ell^{\text{SimPO}}/\partial\pi^{-}$, the analysis is non-trivial. We conclude it as a lemma.
Lemma 1.
As $\pi_{\theta}(a^{-}|x)\to 0$, the limit of the partial derivative regarding $\pi_{\theta}(a^{-}|x)$ in SimPO is:
$\lim_{\pi_{\theta}(a^{-}|x)\to 0}\frac{\partial\ell^{\text{SimPO}}}{\partial\pi_{\theta}(a^{-}|x)}=\begin{cases}0,&\text{if }\beta>|a^{-}|,\\ +\infty,&\text{if }\beta<|a^{-}|,\\ \frac{\beta C}{|a^{-}|},&\text{if }\beta=|a^{-}|,\end{cases}$ (17)
where $C=e^{-\frac{\beta}{|a^{+}|}\log\pi_{\theta}(a^{+}|x)+\gamma}$ is a constant independent of $\pi_{\theta}(a^{-}|x)$.
Proof.
Let $\pi^{-}=\pi_{\theta}(a^{-}|x)$ and $\pi^{+}=\pi_{\theta}(a^{+}|x)$. The partial derivative becomes:
$f(\pi^{-})=\frac{\beta}{|a^{-}|\cdot\pi^{-}}\cdot\sigma(-z),$
where $z=\frac{\beta}{|a^{+}|}\log\pi^{+}-\frac{\beta}{|a^{-}|}\log\pi^{-}-\gamma$.
As $\pi^{-}\to 0$, $\log\pi^{-}\to-\infty$, thus $z\to+\infty$. The sigmoid function approximates to:
$\sigma(-z)\approx e^{-z}=C\cdot(\pi^{-})^{\frac{\beta}{|a^{-}|}},$
where $C=e^{-\frac{\beta}{|a^{+}|}\log\pi^{+}+\gamma}$. By substituting back, we have:
$f(\pi^{-})\approx\frac{\beta C}{|a^{-}|}\cdot(\pi^{-})^{\frac{\beta}{|a^{-}|}-1}.$
Taking the limit as $\pi^{-}\to 0$:
$\lim_{\pi^{-}\to 0}f(\pi^{-})=\begin{cases}0,&\text{if }\frac{\beta}{|a^{-}|}-1>0\ (\beta>|a^{-}|),\\ +\infty,&\text{if }\frac{\beta}{|a^{-}|}-1<0\ (\beta<|a^{-}|),\\ \frac{\beta C}{|a^{-}|},&\text{if }\frac{\beta}{|a^{-}|}-1=0\ (\beta=|a^{-}|).\end{cases}$
Note that in practice, $\beta$ is normally chosen less than $|a^{-}|$, which means the drastic drop in rejected response will happen and 3D-properties still occur.
@@ B.3.3 Simple Preference Optimization (SimPO)
These variants all alleviate the 3D-properties problem of DPO, so the results on mathematical reasoning are partly improved (see Table 9). The invariants we tested do not perform well uniformly on all the tasks. For example, on instruction following tasks like poem generation, vanilla DPO outperforms SLiC and IPO. One hypothesis is their solution forms diverge significantly from the Bradley-Terry model, leading to a loss of generalization in preference learning.
@@ Appendix C Dataset Description
MATH SuperCLUE Poem Slogan UltraFeedBack HH-RLHF
train set 5,826 - 93,269 13,592 170,000 336,820
test set 2,000 1,072 1,000 1,000 - -
@@
Table 6 shows the amount of data in each dataset. MATH, SuperCLUE, UltraFeedBack and HH-rlfh are open-source datasets, while Poem, Slogan are in-house self-built datasets. We provide part of the in-house datasets in the supplementary material to clarify the format and the content. SuperCLUE is only for cross-dataset testing, and COMMON is only used for the comparison between the RM training and DPO training in Section 4.4.